# iWalker: Imperative Visual Planning for Walking Humanoid Robot

### v4

@@ III-B Dynamics-aware Path Planning (iPath)
where we use $[x_{k},y_{k},v_{k},\theta_{k}]^{T}$ as the state vector, representing the robot’s position $x_{k},y_{k}$, velocity $v_{k}$, and orientation $\theta_{k}$; $[\omega_{k},a_{k}]^{T}$ is the control input, consisting of angular velocity $\omega_{k}$ and the acceleration $a_{k}$. To solve this non-linear unicycle dynamics we use an iterative QP solver like iLQR [42] implemented by OSQP [43], and PyPose [44] modules are used to realize BLO. A low $U_{\text{path}}$ indicates that the predicted $\hat{\phi}$ is close to the optimized $\phi^{*}$, and a low $L_{\text{path}}$ means it is easy for MPC to track $\hat{\phi}$ given costs and constraints.
@@
       $U_{\text{ESDF}}\leftarrow U_{\text{ESDF}}+\text{sample}(M_{\text{ESDF}},\hat{s_{x}},\hat{s_{y}})$
@@ III-C Dynamics-aware Mid-level Step Control (iStepper)
$\alpha_{k}$ =$=\begin{cases}1,&\text{if }(k+o)\text{ is even},\\ -1,&\text{if }(k+o)\text{ is odd},\end{cases}$ (14)
@@ III-E End-to-end Training
where $U_{\text{path}}$ exclusively optimizes the iPath module and $U_{\text{step}}$ influences both iPath and iStepper, enabling full differentiability with respect to the vision input. Training is performed with a batch size of 1 using the Adam optimizer. Data augmentation is applied by horizontally flipping both the depth image and the corresponding goal position. Our dataset comprises 1185 frames collected from 12 different indoor scenes using the RealSense camera. Each frameontains monochrome and depth image and the goal position typically set at the end of the sequence. During training, the MPCs are run after the predictions of iPath and iStepper to optimize low-level losses and generate high-level losses. Finally, a single backward pass is performed, ensuring that the entire pipeline benefits from end-to-end gradient flow.
@@ IV-C Quantitative Comparison
We evaluated the performance of our methods and baselines using the following metrics on four recorded test scene sequences: Stair, Office, Corridor, and Corner. onThe Dynamical Feasibility To measure the dynamical feasibility of predicted steps, we measure the low-level costs from unicycle and H-LIP MPC. Intuitively, lower MPC costs indicate less violation of the robot’s dynamics and easier actuatbetter performance Table IV, iWalker, trained iPath and iStepper in an end-to-end manner, eliminates the compound errors and achieves significant improvement, while both iPlanner and iPlanner are unaware of dynamics needed for low-level control and thus have higher cost values.
IV-C1 Collision Safety
@@ Acknowledgements
This work was partially funded by the U.S. Defense Advanced Research Projects Agency (DARPA) grant DARPA-PS-23-13. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of DARPA.

### v5

@@ III-B Dynamics-aware Path Planning (iPath)
where we use $x_{k},y_{k},v_{k},\theta_{k}]^{T}$ as the state vector, representing the robot’s position $x_{k},y_{k}$, velocity $v_{k}$, and orientation $\theta_{k}$; $\omega_{k},a_{k}]^{T}$ is the control input, consisting of angular velocity $\omega_{k}$ and the acceleration $a_{k}$. To solve this non-linear unicycle dynamics we use an iterative QP solver like iLQR [42] implemented by OSQP [43], and PyPose [44] modules are used to realize BLO. A low $U_{\text{path}}$ indicates that the predicted $\hat{\phi}$ is close to the optimized $\phi^{*}$, and a low $L_{\text{path}}$ means it is easy for MPC to track $\hat{\phi}$ given costs and constraints.
@@
[…]
@@ III-C Dynamics-aware Mid-level Step Control (iStepper)
$\alpha_{k}$ $=\begin{cases}1,&\text{if }(k+o)\text{ is even},\\ -1,&\text{if }(k+o)\text{ is odd},\end{cases}$ (14)
@@ III-E End-to-end Training
where $U_{\text{path}}$ exclusively optimizes the iPath module and $U_{\text{step}}$ influences both iPath and iStepper, enabling full differentiability with respect to the vision input. Training is performed with a batch size of 1 using the Adam optimizer. Data augmentation is applied by horizontally flipping both the depth image and the corresponding goal position. Our dataset comprises 1185 frames collected from 12 different indoor scenes using the RealSense camera. Each frame contains monochrome and depth image and the goal position typically set at the end of the sequence. During training, the MPCs are run after the predictions of iPath and iStepper to optimize low-level losses and generate high-level losses. Finally, a single backward pass is performed, ensuring that the entire pipeline benefits from end-to-end gradient flow.
@@ IV-C Quantitative Comparison
We evaluated the performance of our methods and baselines using the following metrics on four recorded test scene sequences: Stair, Office, Corridor, and Corner.
IV-C1 The Dynamical Feasibility
To measure the dynamical feasibility of predicted steps, we measure the low-level costs from unicycle and H-LIP MPC. Intuitively, lower MPC costs indicate less violation of the robot’s dynamics and better performance. As shown in Table IV, iWalker, trained iPath and iStepper in an end-to-end manner, eliminates the compound errors and achieves significant improvement, while both iPlanner and iPlanner are unaware of dynamics needed for low-level control and thus have higher cost values.
IV-C2 Collision Safety
@@ Acknowledgements
This work was supported by the DARPA award HR00112490426. Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the authors and do not necessarily reflect the views of DARPA.